Given that z = 10 - 3i and w = 4 + 2i are complex numbers, find: a) zw b) z/w
write the answer in rectangular form. Help?

Answer by ikleyn(52781)

.
Given that z = 10 - 3i and w = 4 + 2i are complex numbers, find: a) zw b) z/w
write the answer in rectangular form. Help?
~~~~~~~~~~~~~~~~~~~~~~~

zw = (10 - 3i)*(4 + 2i) = 10*4 + 10*2i + (-3i)*4 + (-3i)*(2i) = 

                           40  + 20i   -  12i    + (-6i^2)    =   (use i^2 = -1)

                           40  + (20-12)i        + (-6)*(-1)  = 

                           40  +  8i             + 6          =

                           46  +  8i.

See the lesson
    - Solved problems on arithmetic operations on complex numbers
in this site.